suppose that the width of a certain rectangle is 2 inches more than one fourth of its length. the perimeter is 54 inches. find the length and the width.
Answer by palanisamy(496)

Let the length = x
Given, the width of a certain rectangle is 2 inches more than one fourth of its length.
So, width = x/4 +2
Perimeter is 2*(x+x/4+2) = 54
2x+x/2+4 = 54
(4x+x)/2 = 54-4
5x = 2*50
5x = 100
x = 100/5
x = 20 inches
Therefore the length = 20
width = x/4+2 = 20/4+2 = 5+2 = 7 inches